# Balancing Passenger Transport and Power Distribution: A Distributed Dispatch Policy for Shared Autonomous Electric Vehicles

### v1

@@ I Introduction
Advancements in shared mobility-on-demand services, vehicle automation, and electrification are reshaping the transportation landscape, enabling shared autonomous electric vehicles (SAEVs) to facilitate dynamic interaction between transportation networks and the electric grid [1, 2]. Such vehicles have the potential to reduce electricity demand and voltage fluctuation and improve reliability and resilience in the electric grid if charging and discharging schemes are collaboratively optimized [3, 4, 5, 6, 7]. However, though many studies on vehicle routing have neglected the impact of electric vehicles (EVs) on the grid by assuming infinite power availability [8, 9], research has shown that incorrectly managed EV charging could lead to power quality problems [10]. Appropriate control can negate these concerns and lead to benefits such as peak shaving and voltage and frequency control, particularly when vehicle-to-grid charging technology is employed [11]. Furthermore, integrating EVs with renewable energy sources has shown particular promise in helping align electricity demand and supply curves, increasing economic benefits, and improving resilience [12, 13]. As the SAEV fleet size must be large to serve peak hour demand, the incorporation of constraints related to the energy grid can enable additional uses for these vehicles when they are not needed for transportation.
Beyond day-to-day operations, there are additional resilience benefits when strains are put on the electric grid [14]. Natural disasters can endanger the distribution system, causing equipment failures, blackouts, and even larger scale propagation of failures throughout the network [15]. In the event of electric outages, the extra battery capacity of SAEVs can be leveraged to continue providing power to affected areas, particularly critical facilities such as hospitals and shelters [16, 17, 14]. While some research has examined modifying bus routes to support grid restoration while maintaining transit operations [18], most work has focused on transportable energy storage systems (TESSs), which requires grid operators or government agencies to own and operate a fleet of dedicated vehicles that do not need to transport passengers but can be costly [3]. When SAEVs are used, the benefits to electric consumers must be weighed against the continued needs of transportation passengers who may continue to be reliant on the SAEV fleet after the disaster.
@@
This study introduces a dispatch policy that coordinates the SAEV system with the distribution network. In our framework, the grid operator manages power dispatch while the vehicle dispatcher optimizes vehicle logistics. Unlike previous research, we acknowledge that even during disruptions to the power grid, vehicles will continue to serve passengers. With this in mind, our proposed model ensures that the integrated efforts of the vehicle dispatcher and power grid operator achieve a balanced approach to maintaining essential transportation services, particularly for critical workers, and supporting end-user power needs during system disruptions [19, 20]. This novel cooperative approach allows for joint predictive control and modeling of vehicle and power flows, allowing vehicles to provide services to the electric grid while still maximizing passenger throughput [21, 5, 22, 23]. Furthermore, we advance our methodology by developing a distributed solution approach that enables local decision-making with minimal communication. This approach addresses scalability, stability, privacy, and resilience concerns more effectively. This heuristic, based on the alternating direction method of multipliers (ADMM) [24] achieves near-optimal, dynamic solutions quickly. Our findings from simulations on multiple networks illustrate the ability of the proposed dispatch policy to harmonize the competing demands of transportation and power grids, ultimately highlighting the critical need to integrate operational constraints and objectives from both sectors into the planning model.
@@ II-A Passenger Queuing Constraints
Passegers are tracked based on a queuing model, and vehicle and passenger movements are tracked across the time horizon. We can constrain the dispatch of SAEVs traveling from $q$ to $r$ depending on whether vehicle $v$ is actually parked at node $q$. To do this we define $C_{qs}$ to be the travel time between nodes $q$ and $s$. This travel time is assumed to be constant (not impacted by SAEV dispatch) but could include congestion. Then, vehicle state evolves as:
@@
We consider exogenous demand $d_{qr}(t_{\tau})$ to enter the network at node $r$ with destination $s$ at time $(t_{\tau})$. These passengers form a separate queue $w_{qr}(t_{\tau})$ at each origin for each destination. Queues evolve within the model predictive control framework based on predicted future demand ($\tilde{d}_{qr}(t_{\tau})$). Then, $w_{qr}(t_{\tau})$ evolves as follows:
@@ II-B Vehicle Charging Constraints
where $ep^{v}_{q}(t_{\tau})$ and $eq^{v}_{q}(t_{\tau})$ denote the active and reactive power that any individual vehicle parked at node $q\in N_{R}$ takes from or gives to the grid. As in Singh and Tiwari [25], we assume that EVs are equipped with bi-directional chargers that can inject/absorb reactive power without affecting the state of charge or battery life. Based on the active power taken from the grid, each vehicle is limited in the amount of reactive power that they can take [26, 27]. We can then constrain the reactive power flow based on a linear approximation of the real quadratic constraint (where $\bar{es}_{q}^{v}$ is the maximum allowed real power flow):
@@ II-C Grid Topology Constraints
The LinDistFlow power flow model described in Section II-D assumes a radial network, so the distribution system may need to be reconfigured [18, 31]. The variable $u_{ij}(t_{\tau})\in\{0,1\}$ denotes the line state; $u_{ij}(t_{\tau})=0$ if a line has failed. We also define $s_{ij}(t_{\tau})\in\{0,1\}$ to indicate whether power flow should be allowed to flow. $s_{ij}^{d}$ and $s_{ji}^{r}$ are defined similarly for directed line flow between $(i,j)$ and $(j,i)$ respectively. Since power may only flow in one direction on each line, the topology is constrained by:
@@ II-E Objective Function
In an ideal scenario, the dispatch policy ($\pi^{\star}$) would serve all passenger and electricity demand at every timestep. In this case, the goal should be to minimize either the total electricity consumption or the cost of electricity. However, this is unlikely to be the case in a disrupted network where vehicles may have competing demands from passengers and electric loads. In such a scenario, it makes more sense to maximize the profit or social welfare (which will also minimize electricity consumption/cost where possible). To do this we add up revenues from serving passengers and electric loads and subtract the cost of generating electricity and the cost of excessive charging and discharging of batteries. The objective function is as follows:
@@
In the objective (II-E), the $\mu\geq 0$ functions are pricing functions set by the operator. Cost $\mu_{0}$ is a demand responsive cost that should increase the payments by customers as demand in their queue increases. Cost $\mu_{1}$ represents revenue from passengers once they are dropped at their destination based on distance or average travel times. Cost $\mu_{2}$ represents revenue generated by payments for energy supplied to customers. Cost $\mu_{3}$ is the price of generating energy. Finally, Cost $\mu_{4}$ is the cost of battery degradation from charging and discharging. Since the initial queue lengths $w_{rs}(t)$ are constant with respect to $\pi^{\star}$, the final optimization problem is a mixed integer linear problem (assuming linear cost functions $\mu$).
@@
The objective function proposed in Kang and Levin [21] for shared autonomous vehicle (SAV) dispatch is proven to stabilize demand if it is possible to do so. Note that within this electric vehicle framework the same is true. That proof relies on the first objective term (the demand responsive price) increasing to infinity if queues increase to infinity. This will eventually prioritize passenger service more than providing power to the electric grid and will ensure vehicles serve the longest queues first to maintain stability. For a more extensive discussion of this property and the impacts of electrification on stable dispatch, see Robbennolt [19].
@@ III Distributed Dispatch
In this section, we propose a distributed solution method which is scalable even for long time horizons and large fleet sizes and which also preserves privacy. In the decentralized algorithm, the vehicle dispatcher and power grid operator each solve their own problems, passing power flow estimates at charging stations and iterating until consensus (Figure 2). Section III-A will demonstrate that to solve the dispatch sub-problem, each vehicle will also solve its own lower-level optimization, passing passenger pickup and charging information to the vehicle dispatcher and iterating until convergence (Figure 3).
The distributed solution method proposed is based on ADMM, a well-established distributed optimization technique. Though ADMM cannot guarantee convergence of mixed-integer linear programs (MILPs), it is a useful heuristic since it allows the problem to be decomposed. At the upper level, while the integer variables defining dispatch remain fixed and the grid remains in a single radial configuration, the goal is to determine how much power SAEVs should take from the grid. This problem becomes a linear program (convex) when the integer routing and radial variables are fixed, so ADMM will converge exactly. While vehicles are switching routes, the discontinuities can cause issues with convergence. However, as the number of vehicles grows, these effects diminish, and the approximation improves. At the same time, as the problem grows (especially as the time horizon gets longer and the number of vehicles grows) the centralized problem becomes much more difficult to solve and the computational benefits of each vehicle optimizing its own route become more pronounced.
@@
Many of the constraints in section II are separable by system (roadway or electric network). Constraints (10) and (11) aggregate the charge taken by vehicles from the grid, and link the two problems. These constraints can be relaxed to form a separable problem. In order to create an algorithm to split the decisions of the vehicle dispatcher from the decisions of the power grid operator, each system can create a local copy of the variables $EP_{i}(t_{\tau})$ and $EQ_{i}(t_{\tau})$. This will allow them to optimize their own system and choose the best values of these variables independently. Then, using consensus ADMM [24], a jointly optimal solution can be achieved. Rewrite the centralized problem as:
@@
where $\sigma$ represents the estimate of the aggregate charge that vehicles take from the grid. $\mu$ represents the consensus term that these estimates should converge to. The separable constraints are Equations (II) – (14) for the roadway network and Equations (15) – (34) for the power grid. This problem takes the form of a consensus ADMM problem with two agents (the power grid operator and the vehicle dispatcher) where $EP_{i}(t_{\tau})$ and $EQ_{i}(t_{\tau})$ have been replaced by their local copies $EP^{R}_{i}(t_{\tau})$ and $EQ^{R}_{i}(t_{\tau})$ in equations (10) and (11). In equations (II-D) and (II-D) they have been replaced by local copies $EP^{E}_{i}(t_{\tau})$ and $EQ^{E}_{i}(t_{\tau})$.
@@
The optimization problem (37) can be solved separately by the vehicle dispatcher and power distribution system operator. Each system will determine optimal vehicle movements and power flows (respectively), and both systems will estimate the optimal power flow at charging stations. At each iteration $k$, only the information about predicted power flows at charging stations will be communicated, allowing each system to independently update their own Lagrangian multipliers and continue to the next iteration until convergence.
@@ III-A Vehicle Dispatch Subproblem
Based on the ADMM formulation in equation (37) and (38), we now formulate the vehicle dispatch subproblem (which must be solved at each iteration). The decision variables which relate vehicles to each other are the dispatch decisions to pick up passengers, as well as the decisions about the amount of power to take from the grid (referred to in aggregate as the vector $z^{v}$ for each vehicle). Then, the lower-level vehicle dispatch problem can be written as:
@@
Note that this MILP includes a large number of integer variables, and the size of the problem still increases rapidly as the size of the roadway network increases. Further, this problem must be solved many times before the solution converges to the solution of problem (II-E), and the dispatch problem (II-E) must be solved at every timestep (every 15 seconds – 1 minute). This means the vehicle dispatch subproblem must be solved very quickly. To achieve a sufficiently fast solution we will again propose an ADMM approach, this time distributing subproblems to each vehicle while retaining a the vehicle dispatcher as a centralized coordinator (see figure 3). The inclusion of a term of the form $\sum_{v\in\mathcal{V}}z^{v}_{qr}$ in the objective and constraints makes this problem a sharing problem which can be written in ADMM form by making a local copy of all shared variables [24]. To do this, replace $z^{v}$ in optimization problem (39) with local copies $\eta^{v}$ and then add the consensus constraint $z^{v}-\eta^{v}=0$ $\forall v\in\mathcal{V}$.

### v2

@@ I Introduction
Advancements in shared mobility-on-demand services, vehicle automation, and electrification are reshaping the transportation landscape, enabling shared autonomous electric vehicles (SAEVs) to facilitate dynamic interaction between transportation networks and the electric grid [1, 2]. Such vehicles have the potential to reduce electricity demand and voltage fluctuation and improve reliability and resilience in the electric grid if charging and discharging schemes are collaboratively optimized [3, 4, 5]. However, though many studies on vehicle routing have neglected the impact of electric vehicles (EVs) on the grid by assuming infinite power availability [6, 7], research has shown that incorrectly managed EV charging could lead to power quality problems [8]. Appropriate control can negate these concerns and lead to benefits such as peak shaving and voltage and frequency control, particularly when vehicle-to-grid charging technology is employed [9]. Furthermore, integrating EVs with renewable energy sources has shown particular promise in helping align electricity demand and supply curves, increasing economic benefits, and improving resilience [10]. As the SAEV fleet size must be large to serve peak hour demand, the incorporation of constraints related to the energy grid can enable additional uses for these vehicles when they are not needed for transportation. The integration of SAEVs with power grid operations is motivated by both economic and regulatory factors. Grid operators can compensate fleet operators for power support services, creating new revenue streams while avoiding investments in dedicated TESS fleets. During emergencies, regulatory frameworks may require critical infrastructure providers, including SAEV fleets, to support grid restoration efforts. This cooperation benefits both systems – grid stability ensures SAEV charging availability, while mobile power support helps restore critical loads.
SAEVs can significantly enhance grid resilience through multiple mechanisms. Studies have shown that strategically positioned EVs can provide backup power during outages [11], support voltage regulation and frequency control through vehicle to grid technology [9], enable mobile energy transfer to isolated sections of the grid [3], and facilitate renewable energy integration [12, 13]. The large battery capacity of SAEV fleets makes them particularly valuable for integration with the electric grid [14, 13]. However, beyond day-to-day operations, there are additional resilience benefits when strains are put on the electric grid [15]. Natural disasters can endanger the distribution system, causing equipment failures, blackouts, and even larger scale propagation of failures throughout the network [16]. While existing research has examined using dedicated transportable energy storage systems (TESSs) for grid restoration [3, 17], these approaches require costly dedicated fleets that serve only power needs, ignore the continued need for passenger transportation during disasters, do not address privacy concerns between grid operators and vehicle dispatchers, and lack real-time coordination between power and transportation networks. Furthermore, while studies have explored using transit vehicles for grid support [18], they focus on fixed-route services and fail to address the unique challenges of coordinating large-scale SAEV fleets that must dynamically balance real-time passenger demands with power distribution needs. These gaps highlight the need for a comprehensive framework that can balance competing demands between passenger service and grid support, preserve privacy between stakeholders, enable real-time implementation, and provide provable passenger throughput guarantees.
Shared autonomous vehicle (SAV) dispatch is fundamentally a dial-a-ride problem which is typically assumed to encompass a much larger fleet size and have stochastic demand. Many agent based studies have developed routing heuristics to demonstrate the potential value of these fleets [19, 20, 21]. However, the range of results suggests a need for an optimization framework that can better control vehicles across scenarios [22, 23, 24]. Analytical frameworks used for optimization of the entire fleet can quickly become computationally expensive unless major assumptions are made, particularly when including vehicle rebalancing [25] and EV charging [26, 27, 28]. Though computationally expensive, these analytical frameworks have been shown to outperform other heuristic models both for standard SAV fleets and SAEV fleets which must charge from the grid [29, 26, 27]. Some approaches have developed analytical proofs of maximum throughput which provide additional important service guarantees [30, 4, 31, 25]. Finally, innovative approaches for fleet control based on reinforcement learning have been shown to perform well in simulation [32, 33]. However, these methods can lead to less intuitive results and they do not lend themselves well to an analytical characterization of their properties. In contrast, this work builds heavily on analytical approaches with maximum throughput guarantees, extending these frameworks to the context of a cooperative framework integrated with the power distribution system. This ensures the same stability properties while also enabling additional services and enhanced resilience for disaster recovery.
@@
The contributions of this study are as follows: First, this study introduces a dispatch policy that coordinates the SAEV system with the power distribution network. In our framework, the grid operator manages power dispatch while the vehicle dispatcher optimizes vehicle logistics. Unlike previous research, we acknowledge that even during disruptions to the power grid, vehicles will continue to serve passengers. With this in mind, our proposed model ensures that the integrated efforts of the vehicle dispatcher and power grid operator achieve a balanced approach to maintaining essential transportation services, particularly for critical workers, and supporting end-user power needs during system disruptions [34, 28]. Second, this novel cooperative approach allows for joint predictive control and modeling of vehicle and power flows, allowing vehicles to provide services to the electric grid while still maximizing passenger throughput [30, 4, 31, 25]. This important property ensures that our policy can serve all passenger demand if any policy can, while also providing ancillary services to the grid. Third, we advance our methodology by developing a novel distributed solution approach that enables local decision-making with minimal communication. The hierarchical approach addresses scalability, stability, privacy, and resilience concerns. This heuristic, based on the alternating direction method of multipliers (ADMM) [35] achieves near-optimal, dynamic solutions quickly. Fourth, our findings from simulations on multiple networks illustrate the ability of the proposed dispatch policy to balance the competing demands of transportation and power grids, ultimately highlighting the critical need to integrate operational constraints and objectives from both sectors into the planning model. The case studies also highlight the computational efficiency of the proposed decomposition approach which is crucial for operationalization.
@@ II-A Passenger Queuing Constraints
Passengers are tracked based on a queuing model, and vehicle and passenger movements are tracked across the time horizon. We can constrain the dispatch of SAEVs traveling from $q$ to $r$ depending on whether vehicle $v$ is actually parked at node $q$. To do this we define the exogenous travel time $C_{qs}$ to be the travel time between nodes $q$ and $s$. This travel time is assumed to be constant (not impacted by SAEV dispatch) but could include congestion external to the dispatch problem. Then, vehicle state evolves as:
@@
We consider exogenous demand $d_{qr}(t)$ to enter the network at node $r$ with destination $s$ at time $(t)$. These passengers form a separate queue $w_{qr}(t_{\tau})$ at each origin for each destination. Since the demand is unknown for all $\tau>0$, queues evolve within the model predictive control framework based on predicted future demand ($\tilde{d}_{qr}(t_{\tau})$). Then, $w_{qr}(t_{\tau})$ evolves as follows:
@@ II-B Vehicle Charging Constraints
where $ep^{v}_{q}(t_{\tau})$ and $eq^{v}_{q}(t_{\tau})$ denote the active and reactive power that any individual vehicle parked at node $q\in N_{R}$ takes from or gives to the grid. As in Singh and Tiwari [36], we assume that EVs are equipped with bi-directional chargers that can inject/absorb reactive power without affecting the state of charge or battery life. Based on the active power taken from the grid, each vehicle is limited in the amount of reactive power that they can take [37]. We can then constrain the reactive power flow based on a linear approximation of the real quadratic constraint (where $\bar{es}_{q}^{v}$ is the maximum allowed real power flow):
@@ II-C Grid Topology Constraints
The LinDistFlow power flow model described in Section II-D assumes a radial network, so the distribution system may need to be reconfigured [18, 41]. This problem is coupled with the vehicle dispatch problem since changing power demands at charging stations can affect the pattern of flows in the distribution system, necessitating a different network topology. This coupling is particularly important during service restoration when SAEVs must be strategically positioned to support isolated sections of the grid. The variable $u_{ij}(t_{\tau})\in\{0,1\}$ denotes the line state; $u_{ij}(t_{\tau})=0$ if a line has failed. We also define $s_{ij}(t_{\tau})\in\{0,1\}$ to indicate whether power flow should be allowed to flow. $s_{ij}^{d}$ and $s_{ji}^{r}$ are defined similarly for directed line flow between $(i,j)$ and $(j,i)$ respectively. Since power may only flow in one direction on each line, the topology is constrained by:
@@ II-E Objective Function
In an ideal scenario, the dispatch policy ($\pi^{\star}$) would serve all passenger and electricity demand at every timestep. In this case, the goal should be to minimize either the total electricity consumption or the cost of electricity. However, this is unlikely to be the case in a disrupted network where vehicles may have competing demands from passengers and electric loads. Therefore, in this paper we maximize the profit or social welfare (which will also minimize electricity consumption/cost where possible). To achieve this, we add up revenues from serving passengers and electric loads and subtract the cost of generating electricity and the battery degradation cost of the batteries. The objective function is as follows:
@@
In the objective (II-E), the $\mu\geq 0$ functions are pricing functions set by the operator. The term $\mu_{0}$ is a demand responsive cost that should increase the payments by customers as demand in their queue increases. Term $\mu_{1}$ represents revenue from passengers once they are dropped at their destination based on distance or average travel times. Term $\mu_{2}$ represents revenue generated by payments for energy supplied to customers. Term $\mu_{3}$ is the price of generating energy. Finally, $\mu_{4}$ is the cost of battery degradation from charging and discharging. Since the initial queue lengths $w_{rs}(t)$ are constant with respect to $\pi^{\star}$, the final optimization problem is a mixed-integer linear problem (MILP), assuming linear cost functions $\mu$.
@@
The objective function proposed in Kang and Levin [30] for SAV dispatch is proven to stabilize demand if it is possible to do so. Within this electric vehicle framework the same is true. That proof relies on the first objective term (the demand responsive price) increasing to infinity if queues increase to infinity. This will eventually prioritize passenger service more than providing power to the electric grid and will ensure vehicles serve the longest queues first to maintain stability. For a more extensive discussion of this property and the impacts of electrification on stable dispatch, see Robbennolt [34].
@@ III Distributed Dispatch
In this section, we propose a distributed solution method which is scalable even for long time horizons and large fleet sizes and which also preserves privacy. In the decentralized algorithm, the vehicle dispatcher and power grid operator each solve their own problems, passing power flow estimates at charging stations and iterating until consensus (Figure 2). Section III-A will demonstrate that to solve the dispatch sub-problem, each vehicle will also solve its own lower-level optimization, passing passenger pickup and charging information to the vehicle dispatcher and iterating until convergence (Figure 3). Separating the problem into this hierarchical structure allows privacy to be preserved at multiple levels (i.e. the privacy of each vehicle’s decisions as well as the privacy of the dispatcher and grid operator). In addition, it enables computational effort to be focused on the vehicle dispatch subproblem which is harder to solve. Finally, this decomposition approach allows for large scale parallelization and could potentially allow the time-scales of the vehicle dispatch and power flow optimization problems to be decoupled in the future.
The distributed solution method proposed is based on ADMM, a well-established distributed optimization technique. Though ADMM cannot guarantee convergence of MILPs, it is a useful heuristic since it allows the problem to be decomposed. At the upper level, while the integer variables defining dispatch remain fixed and the grid remains in a single radial configuration, the goal is to determine how much power SAEVs should take from the grid. This problem becomes a linear program (convex) when the integer routing and radial variables are fixed, so ADMM will converge exactly. While vehicles are switching routes, the discontinuities can cause issues with convergence. However, as the number of vehicles grows, these effects diminish, and the approximation improves. At the same time, as the problem grows (especially as the time horizon gets longer and the number of vehicles grows) the centralized problem becomes much more difficult to solve and the computational benefits of each vehicle optimizing its own route become more pronounced.
Although there are other potential solution algorithms for MILPs such as optimal condition decomposition or Benders decomposition, we believe that the proposed heuristic has substantial benefits. In particular, the hierarchical structure of ADMM allows for privacy preservation at multiple levels and could allow future decoupling of the time-scales of the problem [42]. In addition, though other methods might offer stronger convergence guarantees, the extremely large number of binary variables could make their application more challenging and they lack the high parallelization potential of the distributed ADMM architecture. We will demonstrate in the case studies below that the unique problem structure allows for very fast near-optimal solutions using this specially tailored ADMM approach.
@@
Many of the constraints in section II are separable by system (roadway or electric network). However, constraints (10) and (11) aggregate the charge taken by vehicles from the grid at the charging stations and link the vehicle dispatch and power flow problems. In order to create an algorithm to split the decisions of the vehicle dispatcher from the decisions of the power grid operator, each system can create a local copy of the variables $EP_{i}(t_{\tau})$ and $EQ_{i}(t_{\tau})$ which the vehicle dispatcher uses in (10) and (11) and the power grid operator uses in (II-D) and (II-D). Based on this idea, we rewrite the centralized problem as:
@@
where $\sigma$ represents the estimate of the aggregate charge that vehicles take from the grid (a vector of all $EP_{i}(t_{\tau})$ and $EQ_{i}(t_{\tau})$). $\mu$ represents the consensus term that these estimates should converge to. The separable constraints are equations (II) – (14) for the roadway network (R) and equations (15) – (34) for the power grid (E). This problem takes the form of a consensus ADMM problem with two agents (the power grid operator and the vehicle dispatcher) where $EP_{i}(t_{\tau})$ and $EQ_{i}(t_{\tau})$ have been replaced by their local copies $EP^{R}_{i}(t_{\tau})$ and $EQ^{R}_{i}(t_{\tau})$ in equations (10) and (11). In equations (II-D) and (II-D) they have been replaced by local copies $EP^{E}_{i}(t_{\tau})$ and $EQ^{E}_{i}(t_{\tau})$. No other variables are shared between the two agents, so no other equations must be modified.
@@
The mixed-integer quadratic problems (MIQPs) (37) can be solved separately by the vehicle dispatcher and power distribution system operator. Each system will determine optimal vehicle movements and power flows (respectively), and both systems will estimate the optimal power flow at charging stations. At each iteration $k$, only the information about predicted power flows at charging stations will be communicated, allowing each system to independently update their own Lagrangian multipliers and continue to the next iteration until convergence.
@@ III-A Vehicle Dispatch Subproblem
Based on the ADMM formulation in equation (37) and (38), we now formulate the vehicle dispatch subproblem (which must be solved at each iteration). The decision variables which relate vehicles to each other are the dispatch decisions to pick up passengers ($Y$), as well as the decisions about the amount of power to take from the grid ($ep$ and $eq$) and where to charge or discharge ($\hat{\gamma_{ch}}$, and $\hat{\gamma_{dch}}$), (referred to in aggregate as the vector $z^{v}$ for each vehicle). Then, the lower-level vehicle dispatch problem can be written as:
@@
Note that this MILP includes a large number of integer variables, and the size of the problem still increases rapidly as the size of the roadway network increases. Further, this problem must be solved many times before the solution converges to the solution of problem (II-E), and the dispatch problem (II-E) must be solved at every timestep (every 15 seconds – 1 minute). This means the vehicle dispatch subproblem must be solved very quickly. To achieve a sufficiently fast solution we will again propose an ADMM approach, this time distributing subproblems to each vehicle while retaining a the vehicle dispatcher as a centralized coordinator (see figure 3). The inclusion of a term of the form $\sum_{v\in\mathcal{V}}z^{v}_{qr}$ in the objective and constraints makes this problem a sharing problem which can be written in ADMM form by making a local copy of all shared variables [35]. To do this, replace $z^{v}$ in optimization problem (39) with local copies $\eta^{v}$ and then add the consensus constraint $z^{v}-\eta^{v}=0$, $\forall v\in\mathcal{V}$.